# Two-dimensional FrBD friction models for rolling contact

### v4

@@ 1 Introduction
This section extends the FrBD friction model to cover two-dimensional contact cases, where relative rolling and sliding motions may occur in both the longitudinal and lateral directions. The model is derived by mixing the right doses of physical intuition and algebraic dexterity. For this, three main ingredients are needed: a rheological description of a bristle-like element, a local friction model, and a sufficiently convincing mathematical artifice. The first two modelling ingredients are introduced in Sect. 2.1, and then mixed mathematically in Sect. 2.2.
@@ 3.4 Linear rolling contact model for large spin slips
Before moving to Sect. 3.4.1, it is worth commenting on the predictive capabilities of Model 3.4 in relation to those of 3.2 and 3.3. In this context, the first observation concerns the equivalence between the PDEs (52a) and (55a) for $\varphi_{2}(s)\approx 0$, even in the presence of large components $\varphi_{1}(s)$. Essentially, in the context of tyre-road interactions, Model 3.4 can be employed to study the effect of large camber angles and moderate turn spins. Of course, since the bristle deformation $\bm{z}(\bm{x},s)$ may be expected to be small, Eq. (55a) should constitute a sufficiently good approximation to (52a) even when the second spin component is non-negligible ($\varphi_{2}(s)=\varphi_{\psi}(s)$ for tyres and spherical robots). For wheel-rail systems, small values of $\varphi_{2}(s)$ automatically imply small $\varphi_{1}(s)$, and hence the predictive capabilities of Model 3.4 coincide with those of 3.2 for $\varphi_{2}(s)\approx 0$. In fact, Model 3.2 is typically sufficient when it comes to describing wheel-rail rolling contact phenomena.
@@ 4.1.1 Input-to-state stability
The separation condition expressed by Eq. (63) in Lemma 4.1 is reminiscent of the discussion already initiated in Sect. 3.3. As noted earlier, the instantaneous tilting centre $C_{\varphi_{1}}$ (which in this case is fixed) must lie outside the contact area. If this holds, and if one can draw a line through $C_{\varphi_{1}}$ that does not intersect $\mathscr{C}$, then the PDE (59) can be shown to be ISS. In essence, this condition ensures that spinning effects do not dominate the rolling dynamics. Of course, the PDE (47) implicitly assume small spin slips, and therefore the separation condition of Eq. (63) is not formally needed.
@@ Acknowledgements
  • [5] Kalker JJ. Three-Dimensional Elastic Bodies in Rolling Contact. 1st ed. Springer, Dordrecht (1990). Available from: https://doi.org/10.1007/978-94-015-7889-9.
  • [6] Kalker JJ. Rolling Contact Phenomena. In: Jacobson B, Kalker JJ (eds) Rolling Contact Phenomena. International Centre for Mechanical Sciences, vol. 411. Springer, Vienna (2000).
@@
  • [30] Kalker JJ. Transient phenomena in two elastic cylinders rolling over each other with dry friction. J Appl Mech 37(3):677–688 (1970). Available from: https://doi.org/10.1115/1.3408597.
  • [31] Kalker JJ. Transient rolling contact phenomena. ASLE Trans 14(3):177–184 (1971). Available from: https://doi.org/10.1080/05698197108983240.
  • [32] Kalker JJ. On the Rolling Contact of Two Elastic Bodies in the Presence of Dry Friction. PhD thesis, TH Delft (1967).
@@
  • [42] Kalker JJ. Simplified theory of rolling contact. Delft Progress Report, Series C (1973), 1–10.
  • [43] Kalker JJ. Railway wheel and automotive tyre. Veh Syst Dyn 5(15):255–269 (1979).
@@
$\begin{split}\dot{W}(s)&=-\dfrac{1}{2}\iint_{\mathscr{C}}P(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\bm{h}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (88)
@@
$\begin{split}\dot{W}(s)&\leq-\dfrac{1}{2}\oint_{\partial\mathscr{C}}P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\bar{\bm{V}}(\bm{x})\cdot\hat{\bm{n}}_{\partial\mathscr{C}}(\bm{x})\operatorname{d\!}l+\dfrac{1}{2}\iint_{\mathscr{C}}\theta P(\bm{x})N(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}\\ &\qquad+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\bm{h}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0},\end{split}$ (89)
@@
$\begin{split}\dot{W}(s)&\leq-\dfrac{1}{2}\iint_{\mathscr{C}}\theta\eta P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\bm{h}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (90)
@@
$\begin{split}\dot{W}(s)&\leq-\dfrac{1}{2}\iint_{\mathscr{C}}(\theta\eta-2M_{\tilde{\mathbf{\Sigma}}_{\varphi}}-1)P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\dfrac{1}{2}M_{\mathbf{H}}^{2}\iint_{\mathscr{C}}P(\bm{x})\mathinner{\!\left\lVert\bar{\bm{v}}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (92)
@@
$\begin{split}\dot{W}(s)&\leq-\rho W(s)+\mu\mathinner{\!\left\lVert\bar{\bm{v}}(\cdot,\cdot)\right\rVert}_{\infty}^{2},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (93)
@@
$\begin{split}\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)&=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\mathinner{\dfrac{\partial{}\bm{z}(\bm{x},s)}{\partial{s}}}\operatorname{d\!}\bm{x}\\ &=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\Bigl(\tilde{\mathbf{\Sigma}}(\bm{x},s)\bm{z}(\bm{x},s)+\mathbf{H}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)\bar{\bm{v}}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x}\\ &\qquad-\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (96)
@@
$\begin{split}&-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}-\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{1}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\iint_{\mathscr{C}}V_{\textnormal{r}}p(\bm{x})\bar{\bm{v}}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{2}(\bm{x},s)\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S),\end{split}$ (97)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (99)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr)-\dfrac{1}{2}\iint_{\mathscr{C}}\bigl(\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\bigr)\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (100)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\dot{W}\bigl(\bm{z}(\cdot,s)\bigr),\qquad s\in(0,S).\end{split}$ (101)

### v5

@@ 1 Introduction
This section extends the FrBD friction model to cover two-dimensional contact cases, where relative rolling and sliding motions may occur in both the longitudinal and lateral directions. The model is derived by mixing the right doses of physical intuition and algebraic dexterity. For this, three main ingredients are needed: a rheological description of a bristle-like element, a local friction model, and a sufficiently convincing mathematical artifice. The first two modelling ingredients are introduced in Sect. 2.1, and then mixed mathematically in Sect. $2.$
@@ 3.4 Linear rolling contact model for large spin slips
Before moving to Sect. 3.4.1, it is worth commenting on the predictive capabilities of Model 3.4 in relation to those of 3.2 and $3.$ In this context, the first observation concerns the equivalence between the PDEs (52a) and (55a) for $\varphi_{2}(s)\approx 0$, even in the presence of large components $\varphi_{1}(s)$. Essentially, in the context of tyre-road interactions, Model 3.4 can be employed to study the effect of large camber angles and moderate turn spins. Of course, since the bristle deformation $\bm{z}(\bm{x},s)$ may be expected to be small, Eq. (55a) should constitute a sufficiently good approximation to (52a) even when the second spin component is non-negligible ($\varphi_{2}(s)=\varphi_{\psi}(s)$ for tyres and spherical robots). For wheel-rail systems, small values of $\varphi_{2}(s)$ automatically imply small $\varphi_{1}(s)$, and hence the predictive capabilities of Model 3.4 coincide with those of 3.2 for $\varphi_{2}(s)\approx 0$. In fact, Model 3.2 is typically sufficient when it comes to describing wheel-rail rolling contact phenomena.
@@ 4.1.1 Input-to-state stability
The separation condition expressed by Eq. (63) in Lemma 4.1 is reminiscent of the discussion already initiated in Sect. $3.$ As noted earlier, the instantaneous tilting centre $C_{\varphi_{1}}$ (which in this case is fixed) must lie outside the contact area. If this holds, and if one can draw a line through $C_{\varphi_{1}}$ that does not intersect $\mathscr{C}$, then the PDE (59) can be shown to be ISS. In essence, this condition ensures that spinning effects do not dominate the rolling dynamics. Of course, the PDE (47) implicitly assume small spin slips, and therefore the separation condition of Eq. (63) is not formally needed.
@@ Acknowledgements
  • [5] Kalker $J$. Three-Dimensional Elastic Bodies in Rolling Contact. 1st ed. Springer, Dordrecht (1990). Available from: https://doi.org/10.1007/978-94-015-7889-9.
  • [6] Kalker $J$. Rolling Contact Phenomena. In: Jacobson B, Kalker $J$ (eds) Rolling Contact Phenomena. International Centre for Mechanical Sciences, vol. 411. Springer, Vienna (2000).
@@
  • [30] Kalker $J$. Transient phenomena in two elastic cylinders rolling over each other with dry friction. J Appl Mech 37(3):677–688 (1970). Available from: https://doi.org/10.1115/1.3408597.
  • [31] Kalker $J$. Transient rolling contact phenomena. ASLE Trans 14(3):177–184 (1971). Available from: https://doi.org/10.1080/05698197108983240.
  • [32] Kalker $J$. On the Rolling Contact of Two Elastic Bodies in the Presence of Dry Friction. PhD thesis, TH Delft (1967).
@@
  • [42] Kalker $J$. Simplified theory of rolling contact. Delft Progress Report, Series C (1973), 1–10.
  • [43] Kalker $J$. Railway wheel and automotive tyre. Veh Syst Dyn 5(15):255–269 (1979).
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W(s)}{\operatorname{d\!}{s}}}&=-\dfrac{1}{2}\iint_{\mathscr{C}}P(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\bm{h}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (88)
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W(s)}{\operatorname{d\!}{s}}}&\leq-\dfrac{1}{2}\oint_{\partial\mathscr{C}}P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\bar{\bm{V}}(\bm{x})\cdot\hat{\bm{n}}_{\partial\mathscr{C}}(\bm{x})\operatorname{d\!}l+\dfrac{1}{2}\iint_{\mathscr{C}}\theta P(\bm{x})N(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}\\ &\qquad+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\bm{h}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0},\end{split}$ (89)
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W(s)}{\operatorname{d\!}{s}}}&\leq-\dfrac{1}{2}\iint_{\mathscr{C}}\theta\eta P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\mathbf{\Sigma}}_{\varphi}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad+\iint_{\mathscr{C}}P(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\tilde{\bm{h}}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (90)
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W(s)}{\operatorname{d\!}{s}}}&\leq-\dfrac{1}{2}\iint_{\mathscr{C}}(\theta\eta-2M_{\tilde{\mathbf{\Sigma}}_{\varphi}}-1)P(\bm{x})\mathinner{\!\left\lVert\bm{z}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x}+\dfrac{1}{2}M_{\mathbf{H}}^{2}\iint_{\mathscr{C}}P(\bm{x})\mathinner{\!\left\lVert\bar{\bm{v}}(\bm{x},s)\right\rVert}_{2}^{2}\operatorname{d\!}\bm{x},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (92)
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W(s)}{\operatorname{d\!}{s}}}&\leq-\rho W(s)+\mu\mathinner{\!\left\lVert\bar{\bm{v}}(\cdot,\cdot)\right\rVert}_{\infty}^{2},\qquad s\in\mathbb{R}_{>0}.\end{split}$ (93)
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{z}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}}&=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\mathinner{\dfrac{\partial{}\bm{z}(\bm{x},s)}{\partial{s}}}\operatorname{d\!}\bm{x}\\ &=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\Bigl(\tilde{\mathbf{\Sigma}}(\bm{x},s)\bm{z}(\bm{x},s)+\mathbf{H}\bigl(\bar{\bm{v}}(\bm{x},s),s\bigr)\bar{\bm{v}}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x}\\ &\qquad-\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (96)
@@
$\begin{split}&-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}-\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{z}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}}=\iint_{\mathscr{C}}p(\bm{x})\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{1}(\bm{x},s)\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\iint_{\mathscr{C}}V_{\textnormal{r}}p(\bm{x})\bar{\bm{v}}^{\mathrm{T}}(\bm{x},s)\mathbf{Q}_{2}(\bm{x},s)\bar{\bm{v}}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\hskip 17.00024pt+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S),\end{split}$ (97)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{z}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}}+\dfrac{1}{2}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (99)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{z}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}}-\dfrac{1}{2}\iint_{\mathscr{C}}\bigl(\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\bigr)\bm{z}^{\mathrm{T}}(\bm{x},s)\mathbf{\Sigma}_{0}\bm{z}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (100)
@@
$\begin{split}-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}&\geq\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{z}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}},\qquad s\in(0,S).\end{split}$ (101)